# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v2

@@ Appendix B Renormalization effects due to the fields at the X points
The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha={\rm A,B}}F_{\rm X}^{\alpha}$, where
@@
$\begin{split}&C_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}u_{2}^{\alpha}-v_{1}^{\alpha}v_{2}^{\alpha},\quad C_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}u_{4}^{\alpha}-v_{3}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm X}1}^{\alpha}=u_{1}^{\alpha}v_{2}^{\alpha}+v_{2}^{\alpha}u_{1}^{\alpha},\quad S_{{\rm X}2}^{\alpha}=u_{3}^{\alpha}v_{4}^{\alpha}+v_{3}^{\alpha}u_{4}^{\alpha},\\ &C_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}u_{4}^{\alpha}-v_{1}^{\alpha}v_{4}^{\alpha},\quad C_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}u_{3}^{\alpha}-v_{2}^{\alpha}v_{3}^{\alpha},\\ &S_{{\rm Y}1}^{\alpha}=u_{1}^{\alpha}v_{4}^{\alpha}+v_{1}^{\alpha}u_{4}^{\alpha},\quad S_{{\rm Y}2}^{\alpha}=u_{2}^{\alpha}v_{3}^{\alpha}+v_{2}^{\alpha}u_{3}^{\alpha},\\ &C_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}u_{3}^{\alpha}-v_{1}^{\alpha}v_{3}^{\alpha},\quad C_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}u_{4}^{\alpha}-v_{2}^{\alpha}v_{4}^{\alpha},\\ &S_{{\rm Z}1}^{\alpha}=u_{1}^{\alpha}v_{3}^{\alpha}+v_{1}^{\alpha}u_{3}^{\alpha},\quad S_{{\rm Z}2}^{\alpha}=u_{2}^{\alpha}v_{4}^{\alpha}+v_{2}^{\alpha}u_{4}^{\alpha}.\end{split}$ (34)
By minimizing $F_{\rm X}$ with respect to the X points fields, we obtain their stationary values as
@@
$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{\{}\sum_{m>n}\Big{[}u_{0}^{\alpha}\cos(2\theta^{\alpha}_{mn})+v_{0}^{\alpha}\sin(2\theta^{\alpha}_{mn})\Big{]}(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}$
$+6\Big{(}u_{0}^{\alpha}\cos\bar{\theta}^{\alpha}+v_{0}^{\alpha}\sin\bar{\theta}^{\alpha}\Big{)}\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\Bigg{\}},$ (37)
where $\theta_{mn}\equiv\theta_{m}+\theta_{n}$ and $\bar{\theta}^{\alpha}\equiv\theta_{1}^{\alpha}+\theta_{2}^{\alpha}+\theta_{3}^{\alpha}+\theta_{4}^{\alpha}$. Equation (36) renormalizes $c^{\prime}$ and $c^{\prime\prime\prime}$ as
@@ Appendix C Free energy for the orthogonal double-q configuration
[…]
@@
One can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$. The quadratic form of $\delta\phi$ and $\epsilon$ describe the stability of the phase III2q and when one of the eigenvalues vanishes a second-order transition takes place toward the phase II2q.
@@
Rotating $\theta$ leads to the transition to the phase II2q. Assuming $|\theta|\ll 1$ and $\theta_{0}=0$ in Eq. (39), we obtain the quadratic part of $F$ in $\theta$ as
$F=$ $\Big{[}12\lambda_{b}(\phi_{1}^{2}-\phi_{2}^{2})-\lambda_{c}^{(2)}\phi_{0}(\phi_{1}^{2}-2\phi_{2}^{2})-\frac{1728b^{3}}{a_{\rm X}^{2}}\phi^{2}_{1}\phi^{2}_{2}\Big{]}\phi_{0}\theta^{2}.$ (43)
Thus, a second-order transition to the phase II2q takes place when []delimited-[][$\cdots$][ ⋯ ] in Eq. (43) vanishes.

### v3

@@ Appendix B Renormalization effects due to the fields at the X points
The free energy for the X points fields are sublattice diagonal and $F_{\rm X}=\sum_{\alpha=A,B}F_{\rm X}^{\alpha}$, where
@@
$\begin{split}&C_{\ell 1}^{\alpha}=u_{m}^{\alpha}u_{n}^{\alpha}-v_{m}^{\alpha}v_{n}^{\alpha},\quad C_{\ell 2}^{\alpha}=u_{m^{\prime}}^{\alpha}u_{n^{\prime}}^{\alpha}-v_{m^{\prime}}^{\alpha}v_{n^{\prime}}^{\alpha},\\ &S_{\ell 1}^{\alpha}=u_{m}^{\alpha}v_{n}^{\alpha}+v_{m}^{\alpha}u_{n}^{\alpha},\quad S_{\ell 2}^{\alpha}=u_{m^{\prime}}^{\alpha}v_{n^{\prime}}^{\alpha}+v_{m^{\prime}}^{\alpha}u_{n^{\prime}}^{\alpha},\end{split}$ (34)
where $(\ell,m,n)=({\rm X},1,2),\ ({\rm Y},1,4),\ ({\rm Z},1,3)$, and $(\ell,m^{\prime},n^{\prime})=({\rm X},3,4),\ ({\rm Y},2,3),\ ({\rm Z},2,4)$. By minimizing $F_{\rm X}$ with respect to the X points fields, we obtain their stationary values as
@@
$\delta F_{1}^{\alpha}=$ $-\frac{108b^{3}}{a_{\rm X}^{2}}\Bigg{[}\sum_{m>n}\cos(2\theta^{\alpha}_{mn}-\theta_{0}^{\alpha})(\phi^{\alpha}_{m}\phi^{\alpha}_{n})^{2}\phi_{0}^{\alpha}$
$+6\cos(\bar{\theta}^{\alpha}-\bar{\theta}_{\alpha})\phi^{\alpha}_{1}\phi^{\alpha}_{2}\phi^{\alpha}_{3}\phi^{\alpha}_{4}\phi_{0}^{\alpha}\Bigg{]},$ (37)
where $\bm{\phi}_{0}^{\alpha}=(u_{0}^{\alpha},v_{0}^{\alpha})=\phi^{\alpha}_{0}(\cos\theta^{\alpha}_{0},\sin\theta^{\alpha}_{0})$ and $\theta_{mn}\equiv\theta_{m}+\theta_{n}$ and $\bar{\theta}^{\alpha}\equiv\theta_{1}^{\alpha}+\theta_{2}^{\alpha}+\theta_{3}^{\alpha}+\theta_{4}^{\alpha}$. Equation (36) renormalizes $c^{\prime}$ and $c^{\prime\prime\prime}$ as
@@ Appendix C Free energy for the orthogonal double-q configuration
We show the free energy for the orthogonal double-$\bm{q}$ orders with $\bm{\phi}_{2}\perp\bm{\phi}_{1}$: $\bm{\phi}_{1}=\phi_{1}(\cos\theta,\sin\theta),\bm{\phi}_{2}=\phi_{2}(\cos(\theta+\tfrac{\pi}{2}),\sin(\theta+\tfrac{\pi}{2}))=\phi_{2}(-\sin\theta,\cos\theta)$, with the $\Gamma$ point mode $\bm{\phi}_{0}=\phi_{0}(\cos\theta_{0},\sin\theta_{0})$. Here, we assume $\bm{\phi}_{1}=\bm{\phi}_{1}^{A}=\bm{\phi}_{1}^{B}$ and $\bm{\phi}_{2}=\bm{\phi}_{2}^{A}=-\bm{\phi}_{2}^{B}$ as discussed in the main text. The free energy is given as
@@
Although we will not discuss in detail, we can also see that the deviation $\delta\phi=\phi_{1}-\phi_{2}\neq 0$ from the phase III2q accompanies changes in the angle as $\theta=\pi/4+\epsilon$. The quadratic form of $\delta\phi$ and $\epsilon$ describe the stability of the phase III2q and when one of the eigenvalues vanishes a second-order transition takes place toward the phase II2q.
@@
Rotating $\theta$ leads to the transition to the phase II2q. This accompanies also rotation in $\theta_{0}$ accordingly.